OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are 20 and 2500, respectively.

OpenStudy (johnweldon1993):

Well we need to first find the common ratio \[\large -, 20, -, -, 2500 \] Lets modify that a bit until we find the ratio, lets let 20 = a_1 and 2500 = a_4 \[\large a_n = a_1 \times r^{n-1}\] \[\large 2500 = 20 \times r^{4 - 1}\] \[\large 125 = r^3\] \[\large r = 5\] So we found the common ratio to be 5, we can now solve for the first term since we know each term is just multiplied by 5 to make the next term... a_2 = 20...so a_1 = 20/5 = 4 ...so our explicit rule will be \[\large a_n = 4 \times 5^{n - 1}\]

OpenStudy (anonymous):

thank you so much (:

OpenStudy (johnweldon1993):

Anytime :)